OpenStudy (anonymous):

x^2 =-2 I'm having a brain fart. What happens after this?

OpenStudy (anonymous):

or you can add 2 to both sides and factor it

OpenStudy (anonymous):

or just take the sqrt

hartnn (hartnn):

you learnt about imaginary numbers, right ?

OpenStudy (anonymous):

yes I have

hartnn (hartnn):

good, so you know \(\large \sqrt{-1}=i\) right ? what u get if you take square root on both sides of x^2 = -2 ?

OpenStudy (anonymous):

i^2=-1 so, -1 x -2 =2. So the answer is 2?

hartnn (hartnn):

x^2 = -2 \(\sqrt {x^2} = \pm \sqrt {-2} = \pm \sqrt {-1} \sqrt 2 \\x = \pm i \sqrt 2 \)

hartnn (hartnn):

so your 2 values of x are +i sqrt 2 and -i sqrt 2

OpenStudy (anonymous):

ohh Ok Got it! Thanks!

hartnn (hartnn):

welcome ^_^